OpenStudy (anonymous):

Can someone check this please.. The area of a circle is 540 inches squared, what is the area of the sector if its central angle measures 50 degrees. A = 13, 500 inches squared ???

OpenStudy (anonymous):

Area of a sector is proportional to the central angle so, the fraction of the area equals the fraction of the central angles in the same order. \[A/540=50/360\] \[A=540\times50/360\] \[=75\] 75 square inches

OpenStudy (anonymous):

Thank you!!